In the year 2012, a real estate analyst forecasts that a rental apartment building will generate $5.3 million each year in rent over the five years 2013-2017. Cash expenses are expected to be $4.2 million a year. At the end of five years, the building is expected to sell for $12 million. Real estate investors require a 12 percent return on their investments. Apply present value discounting techniques to value the building.
Answer #1
1] Value of the building is the sum of the PV of the
annual net cash inflows and the PV of the sale
value of the building at EOY 5.
2] Net cash inflow per year = 5300000-4200000 = $          11,00,000.00
3] PV of the annual net cash inflow = 1100000*(1.12^5-1)/(0.12*1.12^5) = $          39,65,253.82
Add: PV of the expected sale value at EOY 5 = 12000000/1.12^5 = $          68,09,122.27
Value of the building $       1,07,74,376.09